# From Informal to Formal – Incorporating and Evaluating LLMs on Natural Language Requirements to Verifiable Formal Proofs

### v3

@@ 3.4 Validation Mechanism
Proof Generation Proof Segment Generation ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny ACSL TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL Ave
@@
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 1.43 1.19 6.06 8.33 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 50.00 71.43 11.11
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 22.86 4.86 13.64 11.11 39.68 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 21.43 15.25
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 1.43 1.05 7.58 5.56 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 50.00 57.14 12.57
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 12.86 3.29 12.12 25.00 12.70 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 57.14 18.39
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 50.00 57.14 11.22
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 12.86 4.07 9.85 25.00 26.98 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 57.14 21.19
starcoder-instruct 15B 7.69 2.13 8.47 30.00 28.57 27.14 1.97 6.06 27.78 0.00 27.27 14.29 30.56 42.86 0.00 11.52 27.69 20.00 35.71 28.57 18.27
deepseek-coder-instruct-v1.5 7B 0.00 2.13 6.78 0.00 7.14 4.29 1.73 3.79 5.56 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 7.14 14.29 7.85
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 21.43 7.14 7.32
deepseek-R1 671B 15.38 8.51 25.42 30 35.71 22.86 10.7 21.21 22.22 49.21 45.45 17.86 36.11 28.57 5.88 13.33 40 6.67 42.86 64.29 27.11
Task-wise Average 8.65 10.61 18.44 17.38 43.57
@@
llama3.1-fma 8B 0.00 6.38 8.47 20.00 57.14 41.43 29.98 21.97 25.00 90.48 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 21.43 21.79
llama3.1-ultrachat 8B 0.00 0.00 1.69 0.00 7.14 0.00 0.00 3.79 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 0.00 3.76
llama3.1-ultrachat-fma 8B 0.00 8.51 10.17 30.00 50.00 41.43 35.72 29.55 33.33 95.24 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 35.71 23.75
llama3.1-tulu 8B 0.00 0.00 1.69 0.00 7.14 0.00 0.71 6.06 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 57.14 50.00 8.91
llama3.1-tulu-fma 8B 0.00 4.26 11.86 30.00 35.71 42.86 36.71 27.27 36.11 98.41 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 50.00 25.16
qwen2.5-fma 7B 0.00 4.26 11.86 10.00 71.43 38.57 27.95 27.27 22.22 87.30 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 35.71 21.43 22.25
qwen2.5-coder-fma 7B 0.00 6.38 18.64 20.00 35.71 44.29 36.50 34.09 33.33 98.41 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 50.00 27.31
deepseek-coder-v1.5-fma 7B 0.00 2.13 16.95 0.00 21.43 34.29 25.30 31.06 25.00 84.13 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 21.43 35.71 20.32
Task-wise Average 12.72 $\uparrow$ 33.64 $\uparrow$ 17.92 9.31 22.92
* -fma: fine-tuned with fm-alpaca. * -ultrachat: fine-tuned with UltraChat. * -tulu: fine-tuned with Tulu3.
* -ultrachat-fma: fine-tuned with both UltraChat and fm-alpaca. * -tulu-fma: fine-tuned with both Tulu3 and fm-alpaca.
@@ 4.1 Experiment Setup
Fine-tuning. Instruction fine-tuning Wei et al. (2021); Sanh et al. (2021); Ding et al. (2023); Ivison et al. (2023) aims to improve a model’s ability to effectively respond to human instructions and has shown strong experimental potential in model enhancement. We select llama3.1-8B Meta (2024), qwen2.5-7B Yang et al. (2024a), and deepseek-coder-7B-v1.5  Guo et al. (2024) as base models for fine-tuning. We selected these three models because they have shown promising capability in tasks such as coding, mathematics, and reasoning, and fine-tuning models in their scale is relatively affordable compared with fine-tuning larger scale models. We fine-tuned the three aforementioned base models over three epochs using a learning rate 2e-5, a warm-up ratio of 0.04, a batch size of 512, and a cosine learning rate scheduler.
@@
llama3.1-instruct 8B 0.33 0.32 0.4 0.47 0.34
llama3.1-instruct 70B 0.38 0.31 0.43 0.5 0.37
qwen2.5-instruct 7B 0.34 0.29 0.35 0.41 0.26
qwen2.5-instruct 72B 0.35 0.31 0.36 0.36 0.24
qwen2.5-coder-instruct 7B 0.37 0.34 0.42 0.44 0.31
qwen2.5-coder-instruct 32B 0.34 0.31 0.36 0.41 0.26
starcoder-instruct 15B 0.39 0.36 0.46 0.33 0.37
deepseek-coder-instruct-v1.5 7B 0.42 0.37 0.47 0.55 0.43
deepseek-coder-instruct 33B 0.43 0.36 0.5 0.54 0.46
@@
llama3.1-fma 8B 0.53 0.28 0.58 0.67 0.5
llama3.1-ultrachat 8B 0.44 0.34 0.47 0.57 0.5
llama3.1-ultrachat-fma 8B 0.55 0.42 0.62 0.65 0.68
llama3.1-tulu 8B 0.44 0.38 0.49 0.52 0.47
llama3.1-tulu-fma 8B 0.64 0.41 0.64 0.7 0.76
qwen2.5-fma 7B 0.57 0.24 0.61 0.63 0.57
qwen2.5-coder-fma 7B 0.59 0.31 0.62 0.75 0.75
deepseek-coder-v1.5-fma 7B 0.51 0.36 0.55 0.64 0.46
Average 0.54 (41% $\uparrow$)
@@
Baseline Fine-tuning Datasets: To distinguish whether the capability improvement is simply because more instruction tuning is applied, we also include two commonly used fine-tuning datasets for comparison. We select UltraChat Ding et al. (2023) and Tulu-V3 Lambert et al. (2024) as baseline fine-tuning datasets for their popularity. In particular, UltraChat is a large-scale dataset of instructional conversations that contains 1.5 million high-quality multi-turn dialogues and covers a wide range of topics and instructions. Tulu-v3 Lambert et al. (2024) embraces new data that is either carefully manually curated for quality or generated from GPT models. It is an enhancement of its previous versions Ivison et al. (2023); Wang et al. (2023), focusing more on core skills of knowledge recall, reasoning, mathematics, coding, instruction following, general chat, and safety.
@@
Experiment Environment. The fine-tuning experiment was conducted on 32 Nvidia A100-40G GPUs, while inference was on a single Nvidia A100-80G GPU with vLLM (Kwon et al., 2023).
@@
To understand the current LLMs’ performance in six tasks, we evaluate 8 LLMs against fm-bench with model size ranges from 7B to 72B. The upper part of Table 3 and the upper part of Table 4 show LLMs’ basic performance without fine-tuning.
@@
Task-wise: LLMs perform the best in generating proof from code (Code2Proof), with an average of 43.57% Pass@1, followed by ProofComplete (18.44%) and ProofInfill (17.38%). In contrast, LLMs fall short in generating both the entire formal proof (8.65%) and the proof segments (10.61%). We analyzed the failures and found that syntax errors account for a large proportion, with 12.15% failures caused by syntax errors (Appendix B). The observation echoes the motivation of prior work Wang et al. (2024) and is reasonable due to the grammar difference between most formal specification languages and other programming languages like Python. Regarding requirement analysis, as shown in the upper part of Table 4, the Bleu scores between the ground-truth description and LLM-generated ones range from 0.24 to 0.55. LLM-wise: Without fine-tuning, DeepSeek-R1 achieved the best average (27.11%), followed by qwen2.5-coder-instruct-32B (21.19%).
@@
Model Size – Larger LLMs generally perform better than smaller LLMs. For example, llama3.1-8B only achieved 1.43% in generating TLA+ segments, while llama3.1-70B boosts to 22.86% in the same task. However, there are several exceptions worth noticing, especially for ProofInfill and Code2Proof. For example, llama-3.1-8B achieved 50% in ProofInfill (ACSL), yet the performance drops to 21.43% using the 70B model. Similar observations can be found in qwen2.5-instruct. The decrease in performance is inherently due to the fine-tuning strategy of these instruction models: they are trained to excel in generating rather than filling in the blanks Fried et al. (2022). Also, we conducted a more detailed examination of generated segments and observed that larger LLMs tend to fill in the proof segments that not only look more plausibly correct and well-organized but also include extra content. The additional content, yet, is either redundant, as it repeats information that appears in the subsequent proof, or is incomplete. Promisingly, recent model developers have noticed such conundrums and refined their fine-tuning strategy for fill-in-the-middle tasks Guo et al. (2024).
TLA Coq Lean Dafny ACSL
Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot
LLMs Size Pass@1 Pass@5 Pass@1 Pass@1 Pass@5 Pass@1 Pass@1 Pass@5 Pass@1 Pass@1 Pass@5 Pass@1 Pass@1 Pass@5 Pass@1
llama3.1-instruct 8B 1.43 4.29 11.43 1.19 1.90 10.12 6.06 8.33 9.85 8.33 11.11 19.44 0.00 3.17 38.10
llama3.1-instruct 70B 22.86 27.14 38.57 4.86 6.01 3.46 13.64 17.42 15.91 11.11 13.89 16.67 39.68 52.38 76.19
qwen2.5-instruct 7B 1.43 2.86 5.71 1.05 1.39 12.80 7.58 9.09 9.85 5.56 5.56 13.89 1.59 3.17 66.67
qwen2.5-instruct 72B 12.86 17.14 24.29 3.29 3.80 21.05 12.12 14.39 19.70 25.00 25.00 25.00 12.70 15.87 90.48
qwen2.5-coder-instruct 7B 2.86 5.71 12.86 2.14 2.99 14.30 12.12 13.64 18.94 11.11 11.11 19.44 3.17 4.76 87.30
qwen2.5-coder-instruct 32B 12.86 15.71 30.00 4.07 4.96 18.40 9.85 15.15 13.64 25.00 27.78 30.56 26.98 33.33 96.83
deepseek-coder-instruct-v1.5 7B 4.29 8.57 4.29 1.73 2.82 8.79 3.79 5.30 17.42 5.56 5.56 19.44 0.00 4.76 60.32
deepseek-coder-instruct 33B 2.86 2.86 18.57 1.94 3.23 11.27 4.55 6.82 22.73 11.11 13.89 22.22 3.17 25.40 92.06
Language-wise Average 12.14 6.15 12.00 15.97 34.92
Average 7.68 10.54 18.22 2.53 3.39 12.52 8.71 11.27 16.01 12.85 14.24 20.83 10.01 17.86 75.99

### v4

@@ 3.4 Validation Mechanism
For ACSL specification verification, we report the results under two modes. First, Basic checking. ACSL uses the ‘frama-c’ tool for syntax checking (i.e., , whether the syntax is correct) and other basic ACSL checking. Second, Verification which involves semantics checking (i.e., program behaviors are consistent with the annotated ACSL specifications) and verification towards the property under verification. ‘ACSL-WP’ uses ‘frama-c’ with the WP plugin for formal verification. This mode requires not only the basic checking (i.e., , syntactically correct, semantically consistent with the program), but also that the properties under verification can be proved.
Note that the invalid outputs (e.g., empty response or responses without formal specifications) are considered incorrect directly and will not go through the verification.
@@
ProofGen SegGen ProofComplete ProofInfill Cd2Prf
LLMs Size TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny ACSL ACSL-WP TLA Coq Lean Dafny TLA Coq Lean Dafny ACSL ACSL-WP ACSL ACSL-WP Ave
@@
llama3.1-instruct 8B 0.00 4.26 1.69 0.00 14.29 0.00 1.43 1.19 6.06 8.33 0.00 0.00 0.00 3.57 5.56 28.57 0.00 4.85 7.69 13.33 21.43 0.00 57.14 0.00 7.47
llama3.1-instruct 70B 7.69 6.38 10.17 20.00 7.14 0.00 24.29 4.86 13.64 11.11 39.68 3.17 27.27 3.57 13.89 28.57 0.00 7.27 24.62 13.33 21.43 0.00 21.43 0.00 12.90
qwen2.5-instruct 7B 0.00 2.13 5.08 0.00 21.43 0.00 1.43 1.05 7.58 5.56 1.59 1.59 0.00 13.10 19.44 28.57 0.00 5.45 18.46 13.33 21.43 0.00 14.29 0.00 7.56
qwen2.5-instruct 72B 7.69 6.38 18.64 20.00 21.43 7.14 12.86 3.29 12.12 25.00 12.70 4.76 0.00 11.90 33.33 28.57 0.00 9.70 24.62 26.67 35.71 0.00 57.14 0.00 15.82
qwen2.5-coder-instruct 7B 0.00 2.13 6.78 0.00 0.00 0.00 2.86 2.14 12.12 11.11 3.17 0.00 0.00 4.76 11.11 14.29 0.00 3.64 23.08 20.00 0.00 0.00 0.00 0.00 4.88
qwen2.5-coder-instruct 32B 0.00 6.38 11.86 0.00 21.43 0.00 12.86 4.07 9.85 25.00 26.98 4.76 27.27 5.95 36.11 42.86 2.94 7.88 33.85 20.00 71.43 14.29 57.14 14.29 19.05
starcoder-instruct 15B 7.69 2.13 8.47 30.00 28.57 0.00 27.14 1.97 6.06 27.78 0.00 0.00 27.27 11.90 25.00 42.86 0.00 9.70 24.62 20.00 35.71 0.00 28.57 0.00 15.23
deepseek-coder-instruct 7B 0.00 2.13 8.47 0.00 0.00 0.00 4.29 1.73 3.79 5.56 0.00 0.00 9.09 2.38 13.89 28.57 0.00 3.03 13.85 33.33 0.00 0.00 0.00 0.00 5.42
deepseek-coder-instruct 33B 0.00 0.00 1.69 0.00 0.00 0.00 2.86 1.94 4.55 11.11 3.17 0.00 0.00 3.57 13.89 42.86 0.00 2.42 23.08 6.67 7.14 0.00 7.14 0.00 5.50
deepseek-r1 671B 30.77 8.51 30.51 30.00 35.71 7.14 22.86 10.70 21.21 22.22 49.21 3.17 45.45 15.48 33.33 28.57 5.88 9.70 24.62 6.67 42.86 21.43 64.29 14.29 24.36
Task-wise average 7.70 9.16 18.31 12.62 16.79
@@
llama3.1-fma 8B 0.00 6.38 8.47 20.00 0.00 0.00 38.57 29.58 21.97 25.00 88.89 17.46 36.36 8.33 11.11 28.57 0.00 0.00 7.69 0.00 21.43 0.00 21.43 0.00 16.30
llama3.1-ultrachat 8B 0.00 6.38 3.39 0.00 0.00 0.00 0.00 0.00 3.79 0.00 0.00 0.00 0.00 5.95 13.89 14.29 0.00 4.85 9.23 0.00 14.29 7.14 0.00 0.00 3.47
llama3.1-ultrachat-fma 8B 0.00 8.51 10.17 30.00 0.00 0.00 41.43 35.79 29.55 33.33 95.24 17.46 18.18 8.33 11.11 28.57 2.94 0.00 7.69 0.00 28.57 0.00 35.71 0.00 18.44
llama3.1-tulu 8B 0.00 2.13 1.69 0.00 0.00 0.00 0.00 0.71 6.06 0.00 0.00 0.00 0.00 5.95 8.33 28.57 0.00 1.82 10.77 0.00 7.14 0.00 14.29 0.00 3.64
llama3.1-tulu-fma 8B 0.00 4.26 11.86 30.00 7.14 0.00 42.86 36.64 27.27 36.11 98.41 17.46 18.18 10.71 11.11 42.86 2.94 2.42 6.15 0.00 35.71 0.00 50.00 0.00 20.50
qwen2.5-fma 7B 0.00 4.26 11.86 10.00 0.00 0.00 38.57 27.95 27.27 22.22 87.30 17.46 27.27 9.52 13.89 28.57 0.00 0.00 7.69 0.00 21.43 0.00 14.29 0.00 15.40
qwen2.5-coder-fma 7B 0.00 6.38 18.64 20.00 7.14 0.00 44.29 36.50 34.09 33.33 98.41 17.46 36.36 9.52 16.67 28.57 5.88 1.82 13.85 13.33 42.86 0.00 42.86 0.00 22.00
deepseek-coder-fma 7B 0.00 2.13 16.95 0.00 0.00 0.00 34.29 25.30 31.06 25.00 84.13 15.87 36.36 8.33 19.44 28.57 0.00 1.82 7.69 6.67 14.29 7.14 14.29 7.14 16.10
Task-wise average 6.50 39.82 $\uparrow$ 20.69 $\uparrow$ 7.22 15.48
* -fma: fine-tuned with fm-alpaca. * -ultrachat: fine-tuned with UltraChat. * -tulu: fine-tuned with Tulu3. * -ultrachat-fma: fine-tuned with both UltraChat and fm-alpaca. * -tulu-fma: fine-tuned with both Tulu3 and fm-alpaca.
* ACSL and ACSL-WP: ‘ACSL’ uses ‘frama-c’ tool for syntax checking and other basic ACSL checking; ‘ACSL-WP’ uses ‘frama-c’ with the WP plugin for specification verification.
The task-wise average of finetuned models is calculated excluding the models (i.e., llama3.1-ultrachat and llama3.1-tulu) that are finetuned without fm-alpaca to remove the influence of the average value of the control group models’ results.
@@ 4.1 Experiment Setup
Fine-tuning. Instruction fine-tuning Wei et al. (2021); Sanh et al. (2021); Ding et al. (2023); Ivison et al. (2023) aims to improve a model’s ability to effectively respond to human instructions and has shown strong experimental potential in model enhancement. We select llama3.1-8B-instruct Meta (2024), qwen2.5-7B-instruct Yang et al. (2024a), and deepseek-coder-7B-instruct-v1.5  Guo et al. (2024) as base models for fine-tuning. These three models are chosen due to their promising performance on tasks such as coding, mathematics, and reasoning, and fine-tuning models in their scale is relatively affordable compared with fine-tuning larger scale models. We fine-tuned the three aforementioned models over three epochs using a learning rate 2e-5, a warm-up ratio of 0.04, a batch size of 512, and a cosine learning rate scheduler.
ReqSplitter
@@
llama3.1-instruct 8B 0.33 0.26 0.35 0.36 0.28
llama3.1-instruct 70B 0.33 0.28 0.35 0.41 0.30
qwen2.5-instruct 7B 0.29 0.25 0.32 0.33 0.19
qwen2.5-instruct 72B 0.33 0.28 0.33 0.34 0.22
qwen2.5-coder-instruct 7B 0.29 0.26 0.35 0.42 0.25
qwen2.5-coder-instruct 32B 0.31 0.27 0.33 0.39 0.23
star-coder-instruct 15B 0.21 0.24 0.29 0.30 0.29
deepseek-coder-instruct 7B 0.34 0.30 0.39 0.49 0.36
deepseek-coder-instruct 33B 0.37 0.29 0.47 0.47 0.43
@@
llama3.1-fma 8B 0.33 0.18 0.49 0.49 0.42
llama3.1-ultrachat 8B 0.35 0.30 0.44 0.40 0.46
llama3.1-ultrachat-fma 8B 0.37 0.30 0.56 0.57 0.62
llama3.1-tulu 8B 0.41 0.30 0.43 0.42 0.39
llama3.1-tulu-fma 8B 0.57 0.33 0.57 0.59 0.71
qwen2.5-fma 7B 0.33 0.13 0.53 0.58 0.38
qwen2.5-coder-fma 7B 0.51 0.21 0.57 0.64 0.73
deepseek-coder-fma 7B 0.43 0.24 0.46 0.54 0.43
Average 44.36 (39% $\uparrow$)
@@
Baseline Fine-tuning Datasets: To distinguish whether the capability improvement is simply because more instruction tuning is applied, we also include two commonly used fine-tuning datasets for comparison. We select UltraChat Ding et al. (2023) and Tulu-V3 Lambert et al. (2024) as baseline fine-tuning datasets, and use llama3.1–8B-base Meta (2024) as the base model due to their popularity. In particular, UltraChat is a large-scale dataset of instructional conversations that contains 1.5 million high-quality multi-turn dialogues and covers a wide range of topics and instructions. Tulu-v3 Lambert et al. (2024) embraces new data that is either carefully manu ally curated for quality or generated from GPT models. It is an enhancement of its previous versions Ivison et al. (2023); Wang et al. (2023), focusing more on core skills of knowledge recall, reasoning, mathematics, coding, instruction following, general chat, and safety.
@@
Experiment Environment. The fine-tuning experiment was conducted on 32 Nvidia A800-40G GPUs, while inference was on Nvidia A100-80G GPUs with vLLM (Kwon et al., 2023).
@@
To understand the current LLMs’ performance in six tasks, we evaluate 10 LLMs against fm-bench with model size ranges from 7B to 671B. The upper part of Table 3 and the upper part of Table 4 show LLMs’ basic performance without fine-tuning.
@@
Task-wise: LLMs perform the best in ProofComplete (18.31%) and Code2Proof (16.79%). In contrast, LLMs fall short in generating both the entire formal proof (7.70%) and the proof segments (9.16%). We analyzed the failures and found that syntax errors account for a large proportion, with 12.43% failures caused by syntax errors (See Appendix B). The observation echoes the motivation of prior work Wang et al. (2024) and is reasonable due to the grammar difference between most formal specification languages and other programming languages like Python. Regarding requirement analysis, as shown in the upper part of Table 4, the Bleu scores between the ground-truth description and LLM-generated ones range from 0.19 to 0.49.
LLM-wise: Without fine-tuning, DeepSeek-R1 achieved the best average (24.36%), followed by qwen2.5-coder-instruct-32B (19.05%).
@@
Model Size: Larger LLMs generally perform better than smaller LLMs. For example, llama3.1-8B only achieved 1.43% in generating TLA+ segments, while llama3.1-70B boosts to 24.29% in the same task. However, there are several exceptions worth noticing, especially for ProofInfill and Code2Proof. For example, llama-3.1-8B achieved 57.14% in Code2Proof (ACSL), yet the performance drops to 21.43% using the llama-3.1-70B model. The decrease in performance is inherently due to the fine-tuning strategy of these instruction models: they are trained to excel in generating rather than filling in the blanks Fried et al. (2022). Also, we conducted a more detailed examination of generated segments and observed that larger LLMs tend to fill in the proof segments that not only look more plausibly correct and well-organized but also include extra content. The additional content, yet, is either redundant, as it repeats information that appears in the subsequent proof, or is incomplete. Promisingly, recent model developers have noticed such conundrums and refined their fine-tuning strategy for fill-in-the-middle tasks Guo et al. (2024).
TLA Coq Lean Dafny ACSL ACSL-WP
Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot Zero-shot Few-Shot
LLMs Size P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1 P@1 P@5 P@1
llama3.1-instruct 8B 1.43 4.29 11.43 1.19 1.90 10.12 6.06 8.33 9.85 8.33 11.11 19.44 0.00 3.17 38.10 0.00 0.00 12.70
llama3.1-instruct 70B 24.29 27.14 38.57 4.86 6.01 3.46 13.64 17.42 15.91 11.11 13.89 16.67 39.68 52.38 76.19 3.17 4.76 14.29
qwen2.5-instruct 7B 1.43 2.86 5.71 1.05 1.39 12.80 7.58 9.09 9.85 5.56 5.56 13.89 1.59 3.17 66.67 1.59 3.17 15.87
qwen2.5-instruct 72B 12.86 17.14 24.29 3.29 3.80 21.05 12.12 14.39 19.70 25.00 25.00 25.00 12.70 15.87 90.48 4.76 4.76 15.87
qwen2.5-coder-instruct 7B 2.86 5.71 12.86 2.14 2.99 14.30 12.12 13.64 18.94 11.11 11.11 19.44 3.17 4.76 87.30 0.00 0.00 17.46
qwen2.5-coder-instruct 32B 12.86 15.71 30.00 4.07 4.96 18.40 9.85 15.15 13.64 25.00 27.78 30.56 26.98 33.33 96.83 4.76 4.76 17.46
deepseek-coder-instruct 6.7B 4.29 8.57 4.29 1.73 2.82 8.79 3.79 5.30 17.42 5.56 5.56 19.44 0.00 4.76 60.32 0.00 0.00 9.52
deepseek-coder-instruct 33B 2.86 2.86 18.57 1.94 3.23 11.27 4.55 6.82 22.73 11.11 13.89 22.22 3.17 25.40 92.06 0.00 1.59 15.87
Language-wise Average 12.20 6.15 11.99 15.97 34.92 6.35
Average 7.86 10.54 18.21 2.53 3.39 12.53 8.71 11.27 16.00 12.85 14.24 20.83 10.91 17.86 75.99 1.79 2.38 14.88
@@
* ACSL and ACSL-WP: ‘ACSL’ uses ‘frama-c’ tool for syntax and basic ACSL checking; ‘ACSL-WP’ uses ‘frama-c’ with the WP plugin for specification verification.
@@
Note that though LLMs are generally good at generating syntactically correct and semantically consistent (i.e., program behaviors are consistent with the annotated ACSL specifications) according to the ‘ACSL’ column in Table 3, the specifications are usually insufficient for the verification according to the ‘ACSL-WP’ column in Table 3. The results in the two ACSL columns indicate that there is a large room to improve for generating not only syntactically/semantically correct specifications, but also that the generated specifications need to be sufficient to verify the properties under verification.